OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope -2/5 and y-intercept (0, -3)

OpenStudy (anonymous):

y=-2/5x-3 :)

OpenStudy (anonymous):

did you get this one bella?

OpenStudy (anonymous):

actually i have 4 problems that i am stuck on no i haven't figured it out as yet

OpenStudy (anonymous):

the formula is y=mx+b where m is the slope and b is the y intercept. since the slope is given and the y intercept is -3, then just plug in the values! :)

OpenStudy (anonymous):

okY Phoebeco i get ya

OpenStudy (anonymous):

just remember the formula y = mx + b m= slope ,,,,b= x intercept plug in the point(0,-3) y = (-2/5)x - 3 answer

OpenStudy (anonymous):

wow this is not the answer i got lol

OpenStudy (anonymous):

sure thing, bellarue! :)

OpenStudy (anonymous):

good luck bella...lol

OpenStudy (anonymous):

hey can you help me with one more prob

OpenStudy (anonymous):

k what is it

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

Find the slope and y-intercept of the line in the given equation. y=3/4x - 2

OpenStudy (anonymous):

Mark o you still there

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

y=mx +b y=3/4x - 2 m=3/4 xint= -2

OpenStudy (anonymous):

just use the same formula, y=mx+b, where m is the slope and b is the y intercept, then you'll see that the slope is 3/4 and the y intercept is -2 :)

OpenStudy (anonymous):

always use the formula there..

OpenStudy (anonymous):

you will always use that formula...lol

OpenStudy (anonymous):

so the formula never changes right

OpenStudy (anonymous):

yes it never change ..lol

OpenStudy (anonymous):

i have no idea why i think it has to change

OpenStudy (anonymous):

no it never change...lol

OpenStudy (anonymous):

ok gimmie a sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

do you know of a program that will allow me to do graphs on my Laptop

OpenStudy (anonymous):

http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html try this, it is an online application, though.

OpenStudy (anonymous):

there are software here http://www.wolframalpha.com/input/?i=compute+1+-+3x+%3E%3Dto+10+and+7x-5+%3E+30

OpenStudy (anonymous):

use that theres example there already

OpenStudy (anonymous):

okay thanks a lot you guys are awesome

OpenStudy (anonymous):

yw bella,,,,,,,good luck now

OpenStudy (anonymous):

last question i promise the rest i can finish tomorrow or in a few hours since its almost 430 am here

OpenStudy (anonymous):

bookmark those graphing softwares so that you wont loose them....lol

OpenStudy (anonymous):

okY hold on let me do that now

OpenStudy (anonymous):

oky done

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+-4x%2B5y%3D-20

OpenStudy (anonymous):

click on that bella, that one is one of your prob

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

question on graphing equation

OpenStudy (anonymous):

go ahead.....

OpenStudy (anonymous):

the problem is asking me to Write the slope-intercept form of the equation of the line shown in the graph

OpenStudy (anonymous):

the graph shows these points

OpenStudy (anonymous):

ok yes they are exactly that points..the slope intercept form is always y=mx +b

OpenStudy (anonymous):

-7, 5 and -7,7 my equation would be -7x + 5y=7y+7x

OpenStudy (anonymous):

wait is that points? like this?(-7, 5) and (-7,7 )

OpenStudy (anonymous):

yes atleast thats what am reading on the graph

OpenStudy (anonymous):

which graph? the one im showing you?

OpenStudy (anonymous):

no dear a test sample from my book

OpenStudy (anonymous):

ah..if you see like these (-7, 5) and (-7,7) these are only points in the graph

OpenStudy (anonymous):

ok so i don't have to write and equation

OpenStudy (anonymous):

yes,,..if they required you to write an equation from thay two points,,

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

see this?(-7, 5) and (-7,7) , in my scratch paper they are vertical line if you join those two points.. and on vertical or horizontal line there is NO SLOPE OF THE LINE

OpenStudy (anonymous):

okay am following

OpenStudy (anonymous):

ok veryyyy goood......anything more? any question?

OpenStudy (anonymous):

no tks your awesome and i am off to bed i have to be up in 3hrs

OpenStudy (anonymous):

ok..thennn what states are you in?

OpenStudy (anonymous):

ok good nityy now good luck.bella.....

OpenStudy (anonymous):

thks coz am gonna need it